Assume P=$500-$4Qd and our fixed cost is $1000 and the variable cost is $12 per unit. At what unit is total revenue maximized? How much revenue is generated?

Determine the firm's breakeven points(s) and show graphically showing total revenue and total cost and labeling the breakeven points.
Answer #1

Given

P=500-4Q

Total Revenue=TR=P*Q=(500-4Q)*Q=500Q-4Q2

Marginal Revenue=MR=dTR/dQ=500-8Q

Set MR=0 for revenue maximization

500-8Q=0

Q=500/8=62.50

Revenue is maximized at Q=62.50

P=500-4Q=500-4*62.50=$250

Maximum Revenue=P*Q=250*62.50=$15625

Total Cost=TC=Fixed Cost+Variable Cost=1000+12Q

Set TC=TR to determine break even quantity

1000+12Q=500Q-4Q2

4Q2-488Q+1000=0

Q2-122Q+250=0

-6 + Q= = 62 - 4ac 2a

Q = 122 (-122)2 - 4 * 1 * 250 2*1

Q=(122+117.8304)/2=119.92

and Q=(122-117.8304)/2=2.08

Break even points are Q=2.08 and Q=119.92

Q P=500-4Q TR=P*Q TC=1000+12*Q
0.00 500 0 1000
2.08 492 1025 1025
2.50 490 1225 1030
5.00 480 2400 1060
7.50 470 3525 1090
10.00 460 4600 1120
12.50 450 5625 1150
15.00 440 6600 1180
17.50 430 7525 1210
20.00 420 8400 1240
22.50 410 9225 1270
25.00 400 10000 1300
27.50 390 10725 1330
30.00 380 11400 1360
32.50 370 12025 1390
35.00 360 12600 1420
37.50 350 13125 1450
40.00 340 13600 1480
42.50 330 14025 1510
45.00 320 14400 1540
47.50 310 14725 1570
50.00 300 15000 1600
52.50 290 15225 1630
55.00 280 15400 1660
57.50 270 15525 1690
60.00 260 15600 1720
62.50 250 15625 1750
65.00 240 15600 1780
67.50 230 15525 1810
70.00 220 15400 1840
72.50 210 15225 1870
75.00 200 15000 1900
77.50 190 14725 1930
80.00 180 14400 1960
82.50 170 14025 1990
85.00 160 13600 2020
87.50 150 13125 2050
90.00 140 12600 2080
92.50 130 12025 2110
95.00 120 11400 2140
97.50 110 10725 2170
100.00 100 10000 2200
102.50 90 9225 2230
105.00 80 8400 2260
107.50 70 7525 2290
110.00 60 6600 2320
112.50 50 5625 2350
115.00 40 4600 2380
117.50 30 3525 2410
119.92 20 2439 2439
120.00 20 2400 2440
122.50 10 1225 2470
125.00 0 0 2500

18000 16000 14000 12000 10000 8000 6000 4000 2000 F119.92 72.08 10.00 20.00 40.00 60.00 80.00 100.00 120.00 140.00